# Dynamic Contextual Pricing with Doubly Non-Parametric Random Utility Models

### v1

@@ Abstract
In the evolving landscape of digital commerce, adaptive dynamic pricing strategies are essential for gaining a competitive edge. This paper introduces novel doubly nonparametric random utility models that eschew traditional parametric assumptions used in estimating consumer demand’s mean utility function and noise distribution. Existing nonparametric methods like multi-scale Distributional Nearest Neighbors (DNN and TDNN), initially designed for offline regression, face challenges in dynamic online pricing due to design limitations, such as the indirect observability of utility-related variables and the absence of uniform convergence guarantees. We address these challenges with innovative population equations that facilitate nonparametric estimation within decision-making frameworks and establish new analytical results on the uniform convergence rates of DNN and TDNN, enhancing their applicability in dynamic environments.
@@ 1.1 Doubly Nonparametric Random Utility Models
The proposed doubly nonparametric random utility model are characterized by (1) and (3) without any parametric characterizations on $\mathring{\mu}(\boldsymbol{x}_{t})$ or $F(\cdot)$. It operates within a broad economic framework of random utility models. In such settings, products are sold individually, with each sale resulting in a binary outcome indicating success or failure. At each decision point $t\in[T]$, a $d$-dimensional contextual covariate $\boldsymbol{x}_{t}$ from $\mathbb{R}^{d}$ is independently and identically drawn from a fixed, but a priori unknown, distribution. The expected market value of the product at time $t$ is determined by the observed contextual covariate $\boldsymbol{x}_{t}$. A general random utility model for the market value of the product is given by
@@
where $\mathring{\mu}(\cdot)$ is the unknown function of the mean utility, and $\varepsilon_{t}$ are i.i.d. noises following an unknown distribution $F(\cdot)$ with $\mathbb{E}[\varepsilon_{t}]=0$. Given a posted price of $p_{t}$ for the product at time $t$, we observe $y_{t}:=\mathbbm{1}(v_{t}\geq p_{t})$ that indicates whether the a sale occurs ($y_{1}=1$) or not ($y_{1}=0$). The model is equivalent to the probabilistic model:
@@
The doubly nonparametric contextual pricing works under a general economic framework of random utility models, the settings of which fundamentally differ from models that directly that directly model the revenue function. For example, Chen and Gallego (2021) directly models the unknown expected revenue as $\operatorname{rev}_{t}(p_{t})=p_{t}\cdot d(\boldsymbol{x}_{t},p_{t})$ with $d(\boldsymbol{x}_{t},p_{t})\in[0,1]$ representing the personalized demand function. Bu et al. (2022) further assumes that $d(\boldsymbol{x}_{t},p_{t})$ separable in terms of $p_{t}$ and $\boldsymbol{x}_{t}$. In contrast, we work under the random utility model; the unknown expected revenue is given by (3) where it can be viewed as $d(\boldsymbol{x}_{t},p_{t})=1-F(p_{t}-\mathring{\mu}(\boldsymbol{x}_{t}))$. This finer granularity in demand decomposition allows us to separately analyze the mean utility function and noise distribution, thus improving theoretical guarantees. Furthermore, our analysis highlights a crucial trade-off between the dimensionality and smoothness of $\mathring{\mu}(\cdot)$ and $F(\cdot)$, offering practitioners a more nuanced control over pricing strategies.
@@ 1.2 Novelty and Contributions
This paper distinct from existing literature by adopting a fully nonparametric approach in modeling both the mean utility function, $\mathring{\mu}(\cdot)$, and the noise distribution, $F(\cdot)$, in the context of dynamic pricing with random utility. This approach eliminates the risk of model misspecification and enhances the flexibility and accuracy of our pricing strategy, which is particularly beneficial in complex and evolving market environments. We employ state-of-the-art nonparametric methods, notably the Distributional Nearest Neighbors (DNN) and the two-scale distributed nearest neighbor (TDNN) techniques, which are known for their simplicity in implementation and minimal assumptions. These estimators not only provide the flexibility needed in dynamic pricing but also achieve faster convergence rates, improving as we scale up the model’s complexity and data fidelity.
@@
However, existing DNN and TDNN methods, traditionally used for offline regression, are not directly applicable to the dynamic and online nature of pricing decisions due to their design limitations. Key challenges include the indirect observability of utility-related variables and the absence of robust convergence guarantees in such a dynamic context. To overcome these limitations, we employed two novel population equations, (5) and (10), which enable the nonparametric estimation of both $\mathring{\mu}(\cdot)$ and $F(\cdot)$ using only binary response data $y_{t}$. This innovative approach bypasses the need for direct measurements of variables like $v_{t}$ or $\epsilon_{t}$, enhancing the robustness and accuracy of our estimations, and thereby substantially reducing the regret associated with pricing decisions.
@@
Our theoretical developments include a new uniform convergence rate in Theorem 5 for the DNN and TDNN estimators. Employing the elegant U-Statistic representation of DNN and TDNN, Theorem 5 establishes that the DNN and TDNN estimators achieve the optimal rate of uniform convergence for nonparametric regression under the second-order smoothness condition and the fourth-order smoothness condition, respectively (Stone, 1982). Moreover, theoretical analyses uncover new insights into the interdependencies between the estimation accuracy $\mathring{\mu}(\cdot)$ and $F(\cdot)$, and how these relationships influence the performance of the overall pricing strategy in the doubly nonparametric setting. Specifically, we demonstrate that the regret bounds for our doubly nonparametric dynamic pricing model depend intricately on the trade-offs between the dimensionality of contextual variables ($d$) and the smoothness properties of the noise distribution ($m$). To simply summarize, when $1\leq d<\frac{6}{m-1}$, the regret is bounded by $\widetilde{\Theta}(T^{\frac{2m+1}{4m-1}})$; and when $d>\frac{6}{m-1}$, the regret is bounded by $\widetilde{\Theta}(T^{\frac{d+4}{d+8}})$, ignoring the logarithmic terms. Regret bounds under both settings improve upon existing nonparametric models (Chen and Gallego, 2021; Bu et al., 2022), offering finer control and better performance in practical settings.
@@ 1.3 Literature and Organization
Contextual dynamic pricing under random choice models. Recently, there has been a growing body of literature on context-based dynamic pricing, where product and customer features are taken into account when modeling the demand curve or market evaluation. We refer the reader to recent surveys (Den Boer, 2015; Ban and Keskin, 2021; Chen and Gallego, 2022; Chen et al., 2022; Chen and Hu, 2023) for a comprehensive review on this stream of literature. The most relevant literature to the present paper is those that model the demand by a binary random choice model characterized by (1) and (2). This group of literature can be categorized by the assumptions on the two unknown functions, namely the mean utility $\mathring{\mu}(\boldsymbol{x}_{t})$ and the CDF of the noise $F(\cdot)$. Assuming the CDF $F(\cdot)$ is a known priori, Javanmard and Nazerzadeh (2019); Cohen et al. (2020); Xu and Wang (2021) all impose a linear parametric model on the expected market value of the product at time $t$, i.e. $\mathring{\mu}(\boldsymbol{x}_{t})=\boldsymbol{\alpha}^{\top}\boldsymbol{x}_{t}$ in (1), where $\boldsymbol{\alpha}^{\top}$ is some unknown parameter, and aim to estimate the mean utility while optimizing the revenue. Alternatively, Shah et al. (2019); Fan et al. (2022); Luo et al. (2024) allow the noise distribution, i.e. the CDF $F(\cdot)$ in (2), to be unknown. (Fan et al., 2022; Luo et al., 2024) assume an underlying linear mean utility $\mathring{\mu}(\boldsymbol{x}_{t})$ and estimate $F$ by kernel estimation and by the discretization method (Kleinberg and Leighton, 2003), respectively. Shah et al. (2019) assumes the market value $v_{t}=\exp(\theta^{T}\boldsymbol{x}_{t}+\varepsilon_{t})$ where $\varepsilon_{t}$ has unknown distribution and proposes the DEEP-C algorithm to optimize the regret. However, they all work under the settings where the mean utility is parametrically organized.
@@
A stream of literature in contextual dynamic pricing models the demand function $d(\boldsymbol{x},p)$ or the revenue function $p\cdot d(\boldsymbol{x},p)$ directly (Qiang and Bayati, 2016; Ban and Keskin, 2021; Chen and Gallego, 2021; Wang et al., 2021; Bu et al., 2022). This stream of work can be categorized into contextual bandits with a specialized reward function $f(\boldsymbol{x}_{t},p)=p\cdot d(\boldsymbol{x},p)$ and continuous scalar action $p$. Under the general bandit setting, Slivkins (2011) studies nonparametric bandits assuming that $f(\boldsymbol{x}_{t},p)$ is Lipschitz and establishes the optimal regret being $\widetilde{\Theta}(T^{\frac{d+2}{d+3}})$ when $p$ is a scalar [^1] Slivkins (2011) works with a general setting where an action $p$ can be a vector.. Mao et al. (2018) exploits a special structure where the demand function is modeled as a strictly binary feedback $\mathbf{1}\{p\leq\mathring{\mu}(\boldsymbol{x}_{t})\}$, where $\mathring{\mu}(\boldsymbol{x}_{t})$ is Lipschitz, and develops the regret bound of $\widetilde{\Theta}(T^{\frac{d}{d+1}})$. Guan and Jiang (2018) applies $K$-Nearest Neighbour to nonparametric stochastic contextual bandits whose rate is not as good $\widetilde{\Theta}(T^{\frac{d+1}{d+2}})$. Under the dynamic pricing setting, Chen and Gallego (2021) estimate $f(\boldsymbol{x}_{t},p)=p\cdot d(\boldsymbol{x},p)$ directly. However, they do not impose any functional structure that is specific to the dynamic pricing problem. Assuming that the revenue function is Lipschitz continuous and locally concave, the authors prove that the optimal regret is $\widetilde{\Theta}(T^{\frac{d+2}{d+4}})$ where $d$ is the dimension of $\boldsymbol{x}$. Nambiar et al. (2019) study the demand function $bp+g(x)$ motivated by model misspecification and propose a random price shock (RPS) algorithm to estimate price elasticity. They establish the optimal regret compared with those who know the best linear approximation to $g(x)$ is of order $\widetilde{\Theta}(\sqrt{T})$. Bu et al. (2022) further explore the separable structure of the demand function in the sense that $d(\boldsymbol{x},p)=a(\boldsymbol{x})+b(p)$ where both function $a(\cdot)$ and $b(\cdot)$ are unknown. This paper adopts the random utility model for demand characterization and is very different from these works in terms of aggregate demand modeling, algorithm development, and regret analysis.
@@
Personalised dynamic pricing with related topics. There has been a wide range of literature studying personalised dynamic pricing with additional interesting topics including inventory constraints, fairness, privacy protection and more. We refer the readers to some recent works that could potentially incorporate our method in future research. In terms of joint pricing and inventory control problem, Chen et al. (2021) develop learning algorithm with censored demand and approximates $f(d_{t},p_{t},y_{t})$ nonparametrically, where $y_{t}$ represents inventory levels. Feng and Zhu (2023) further explores the impact of price protection guarantee, where the customers who purchased a product are endowed the right to receive a refund from the seller if he lowers the price during the price protection period. Li and Zheng (2023) take the inventory constraints into consideration where the inventory level B is given at the beginning and cannot be replenished. All those studies develop learning algorithm to estimate the underlying unknown demand functions, where our method could be seamlessly integrated. In terms of fairness, Maestre et al. (2018) considers the fairness in price, they introduce Jain’s index to measures the fairness in terms of the average price across groups, while Cohen et al. (2021) estimate unknown demand function under both price fairness and time fairness.
@@
Additional constraints or addition conditions of related interest in dynamic pricing are explored in the following works. Chen et al. (2021) adopt the fundamental framework of differential privacy when implementing demand learning, and develop privacy-preserving dynamic pricing, which tries to maximize the retailer revenue while avoiding leakage of individual customer’s information. Keskin and Zeevi (2017) adopt a linear underlying demand model but assumes the unknown parameters can change over time. van den Boer and Keskin (2022) take reference-price effects into account and learn demand uncertainty by adding the reference-price updating mechanism with unknown parameters. Bastani et al. (2021) leverages a potentially shared structure of the demand function cross multiple products and conduct a meta dynamic pricing algorithm to make pricing more efficiently.
@@ 2 Doubly Non-Parametric Learning and Decision
We propose an Algorithm 1 which outputs the empirical optimal policy for contextual dynamic pricing and balances the exploration-exploitation trade-off. The horizon of $T$ periods is divided into episodes with increasing length. Specifically, the $k$th episode has length $n_{k}=2^{k-1}n_{0}$, where $k\geq 1$ and $n_{0}$ is the minimal episode length. For each $k\geq 1$, the $k$th episode is further divided into exploration and commit (or exploitation) phases, which are denoted by ${\cal T}_{k,exp}$ and ${\cal T}_{k,com}$, respectively. In the exploration phase ${\cal T}_{k,exp}$, we collect sample $(x_{t},By_{t})$ through randomly setting price $p_{t}\stackrel{{\scriptstyle d}}{{\sim}}\operatorname{Unif}(0,B)$ and then estimate both $\mathring{\mu}(\cdot)$ or $F(\cdot)$ nonparametrically using the collected sample $\{(x_{t},By_{t})\}_{t\in{\cal T}_{k,exp}}$. In the exploitation phase ${\cal T}_{k,com}$, we proceed to formulate a data-driven optimal policy leveraging the estimated mean utility and noise distribution learned in the exploration phase. The sizes of the exploration phase of uniform random pricing and the exploitation phase of greedy pricing based on estimation can be determined in theoretical analysis to minimize the total regret. The subsequent subsections introduce our proposed approach for the doubly non-parametric estimation of $\mathring{\mu}(\cdot)$ and $F(\cdot)$.
@@
11      Using samples $\{(\boldsymbol{x}_{t},By_{t})\}_{t\in{\cal T}_{k,exp}}$ to fit $\widehat{\mu}_{k}(\boldsymbol{x}_{t})$ for $t\in{\cal T}_{k,exp}$ according to DNN or TDNN (9).
@@
17             Predict $\widehat{\mu}_{k}(\boldsymbol{x}_{t})$ according to DNN or TDNN (9).

### v2

@@ Abstract
In the evolving landscape of digital commerce, adaptive dynamic pricing strategies are essential for gaining a competitive edge. This paper introduces novel doubly nonparametric random utility models that eschew traditional parametric assumptions used in estimating consumer demand’s mean utility function and noise distribution. Existing nonparametric methods like multi-scale Distributional Nearest Neighbors (DNN and TDNN), initially designed for offline regression, face challenges in dynamic online pricing due to design limitations, such as the indirect observability of utility-related variables and the absence of uniform convergence guarantees. We address these challenges with new population equations that facilitate nonparametric estimation within decision-making frameworks and establish the first analytical results on the uniform convergence rates of DNN and TDNN, enhancing their applicability in dynamic environments.
@@ 1.1 Doubly Nonparametric Random Utility Models
The proposed doubly nonparametric random utility model are characterized by (1) and (2) without any parametric characterizations on $\mathring{\mu}(\boldsymbol{x}_{t})$ or $F(\cdot)$. It operates within a broad economic framework of random utility models. In such settings, products are sold individually, with each sale resulting in a binary outcome indicating success or failure. At each decision point $t\in[T]$, a $d$-dimensional contextual covariate $\boldsymbol{x}_{t}$ from $\mathbb{R}^{d}$ is independently and identically drawn from a fixed, but a priori unknown, distribution. The expected market value of the product at time $t$ is determined by the observed contextual covariate $\boldsymbol{x}_{t}$. A general random utility model for the market value $v_{t}$ of the product is given by
@@
where $\mathring{\mu}(\cdot)$ is the unknown function of the mean utility, and $\varepsilon_{t}$ are i.i.d. noises following an unknown distribution $F(\cdot)$ with $\mathbb{E}[\varepsilon_{t}]=0$. We assume that the market value $v_{t}$ is within an interval $[B_{v},B-B_{v}]$ with constants $0<B_{v}<B$. Given a posted price of $p_{t}$ for the product at time $t$, we observe $y_{t}:=\mathbbm{1}(v_{t}\geq p_{t})$ that indicates whether the a sale occurs ($y_{1}=1$) or not ($y_{1}=0$). The model is equivalent to the probabilistic model:
@@
The doubly nonparametric contextual pricing works under a general economic framework of random utility models, the settings of which fundamentally differ from models that directly model the revenue function. For example, Chen and Gallego (2021) directly models the unknown expected revenue as $\operatorname{rev}_{t}(p_{t})=p_{t}\cdot d(\boldsymbol{x}_{t},p_{t})$ with $d(\boldsymbol{x}_{t},p_{t})\in[0,1]$ representing the personalized demand function. Bu et al. (2022) further assumes that $d(\boldsymbol{x}_{t},p_{t})$ separable in terms of $p_{t}$ and $\boldsymbol{x}_{t}$. In contrast, we work under the random utility model; the unknown expected revenue is given by (3) where it can be viewed as $d(\boldsymbol{x}_{t},p_{t})=1-F(p_{t}-\mathring{\mu}(\boldsymbol{x}_{t}))$. This finer granularity in demand decomposition allows us to separately analyze the mean utility function and noise distribution, thus improving theoretical guarantees. Furthermore, our analysis highlights a crucial trade-off between the dimensionality and smoothness of $\mathring{\mu}(\cdot)$ and $F(\cdot)$, offering practitioners a more nuanced control over pricing strategies.
@@ 1.2 Novelty and Contributions
This paper distincts from existing literature by adopting a fully nonparametric approach to model both the mean utility function, $\mathring{\mu}(\cdot)$, and the noise distribution, $F(\cdot)$, in the context of dynamic pricing with random utility. This approach eliminates the risk of model misspecification and enhances the flexibility and accuracy of our pricing strategy, which is particularly beneficial in complex and evolving market environments. We employ state-of-the-art nonparametric methods, notably the Distributional Nearest Neighbors (DNN) and the two-scale distributed nearest neighbor (TDNN) techniques, which are known for their simplicity in implementation and minimal assumptions. These estimators not only provide the flexibility needed in dynamic pricing but also achieve optimal convergence rates, improving as we scale up the model’s complexity and data fidelity.
@@
However, existing DNN and TDNN methods, traditionally used for offline regression, are not directly applicable to the dynamic and online nature of pricing decisions due to their design limitations. Key challenges include the indirect observability of utility-related variables and the absence of robust convergence guarantees in such a dynamic context. To overcome these limitations, we employed two new population equations, (5) and (11), which enable the nonparametric estimation of both $\mathring{\mu}(\cdot)$ and $F(\cdot)$ using only binary response data $y_{t}$. This innovative approach bypasses the need for direct measurements of variables like $v_{t}$ or $\epsilon_{t}$, enhancing the robustness and accuracy of our estimations, and thereby substantially reducing the regret associated with pricing decisions.
@@
Our theoretical developments include a new uniform convergence rate in Theorem 5 for the DNN and TDNN estimators. Employing the elegant U-Statistic representation of DNN and TDNN, Theorem 5 establishes that the DNN and TDNN estimators achieve the optimal rate of uniform convergence for nonparametric regression under the second-order smoothness condition and the fourth-order smoothness condition, respectively (Stone, 1982). Moreover, theoretical analyses uncover new insights into the interdependencies between the estimation accuracy for $\mathring{\mu}(\cdot)$ and $F(\cdot)$, and how these relationships influence the performance of the overall pricing strategy in the doubly nonparametric setting. Specifically, we demonstrate that the regret bounds for our doubly nonparametric dynamic pricing model depend intricately on the trade-offs between the dimensionality of contextual variables ($d$) and the smoothness properties of the noise distribution ($m$). To simply summarize, when $1\leq d\leq\frac{6}{m-1}$, the regret is bounded by $\widetilde{{\cal O}}(T^{\frac{2m+1}{4m-1}})$; and when $d>\frac{6}{m-1}$, the regret is bounded by $\widetilde{{\cal O}}(T^{\frac{d+4}{d+8}})$, ignoring the logarithmic terms. Regret bounds under both settings improve upon existing nonparametric models (Chen and Gallego, 2021; Bu et al., 2022), offering finer control and better performance in practical settings.
@@ 1.3 Literature and Organization
Contextual dynamic pricing under random choice models. Recently, there has been a growing body of literature on context-based dynamic pricing, where product and customer features are taken into account when modeling the demand curve or market evaluation. We refer the reader to recent surveys (Den Boer, 2015; Ban and Keskin, 2021; Chen and Gallego, 2022; Chen et al., 2022; Chen and Hu, 2023) for a comprehensive review on this stream of literature. The most relevant literature to the present paper is those that model the demand by a binary random choice model characterized by (1) and (2). This group of literature can be categorized by the assumptions on the two unknown functions, namely the mean utility $\mathring{\mu}(\boldsymbol{x}_{t})$ and the CDF of the noise $F(\cdot)$. Assuming the CDF $F(\cdot)$ is a known priori, Javanmard and Nazerzadeh (2019); Cohen et al. (2020); Xu and Wang (2021) all impose a linear parametric model on the expected market value of the product at time $t$, i.e. $\mathring{\mu}(\boldsymbol{x}_{t})=\boldsymbol{\alpha}^{\top}\boldsymbol{x}_{t}$ in (1), where $\boldsymbol{\alpha}$ is some unknown parameter, and aim to estimate the mean utility while optimizing the revenue. Alternatively, Shah et al. (2019); Fan et al. (2022); Luo et al. (2024) allow the noise distribution, i.e. the CDF $F(\cdot)$ in (2), to be unknown. (Fan et al., 2022; Luo et al., 2024) assume an underlying linear mean utility $\mathring{\mu}(\boldsymbol{x}_{t})$ and estimate $F$ by kernel estimation and by the discretization method (Kleinberg and Leighton, 2003), respectively. While Shah et al. (2019) assume the market value $v_{t}=\exp(\boldsymbol{\alpha}^{T}\boldsymbol{x}_{t}+\varepsilon_{t})$, where $\varepsilon_{t}$ has unknown distribution and propose the DEEP-C algorithm to optimize the regret. However, they all work under the settings where the mean utility is parametrically organized.
@@
A stream of literature in contextual dynamic pricing models the demand function $d(\boldsymbol{x},p)$ or the revenue function $p\cdot d(\boldsymbol{x},p)$ directly (Qiang and Bayati, 2016; Ban and Keskin, 2021; Chen and Gallego, 2021; Wang et al., 2021; Bu et al., 2022). This stream of work can be categorized into contextual bandits with a specialized reward function $f(\boldsymbol{x}_{t},p)=p\cdot d(\boldsymbol{x},p)$ and continuous scalar action $p$. Under the general bandit setting, Slivkins (2011) studies nonparametric bandits assuming that $f(\boldsymbol{x}_{t},p)$ is Lipschitz and establishes the optimal regret being $\widetilde{{\cal O}}(T^{\frac{d+2}{d+3}})$ when $p$ is a scalar [^1] Slivkins (2011) works with a general setting where an action $p$ can be a vector.. Mao et al. (2018) exploit a special structure where the demand function is modeled as a strictly binary feedback $\mathbf{1}\{p\leq\mathring{\mu}(\boldsymbol{x}_{t})\}$, where $\mathring{\mu}(\boldsymbol{x}_{t})$ is Lipschitz, and develops the regret bound of $\widetilde{{\cal O}}(T^{\frac{d}{d+1}})$. Guan and Jiang (2018) apply $K$-Nearest Neighbour to nonparametric stochastic contextual bandits attaining a regret of $\widetilde{{\cal O}}(T^{\frac{d+1}{d+2}})$. Under the dynamic pricing setting, Chen and Gallego (2021) estimate $f(\boldsymbol{x}_{t},p)=p\cdot d(\boldsymbol{x},p)$ directly. However, they do not impose any functional structure that is specific to the dynamic pricing problem. Assuming that the revenue function is Lipschitz continuous and locally concave, the authors prove that the optimal regret is $\widetilde{{\cal O}}(T^{\frac{d+2}{d+4}})$ where $d$ is the dimension of $\boldsymbol{x}$. Nambiar et al. (2019) study the demand function $bp+g(x)$ motivated by model misspecification and propose a random price shock (RPS) algorithm to estimate price elasticity. They establish the optimal regret of order $\widetilde{{\cal O}}(\sqrt{T})$ compared to a clairvoyant who knows the best linear approximation to $g(x)$. Bu et al. (2022) further explore the separable structure of the demand function in the sense that $d(\boldsymbol{x},p)=a(\boldsymbol{x})+b(p)$ where both function $a(\cdot)$ and $b(\cdot)$ are unknown. This paper adopts the random utility model for demand characterization and is very different from these works in terms of aggregate demand modeling, algorithm development, and regret analysis.
@@
Contextual dynamic pricing with related topics. There has been a wide range of literature studying contextual dynamic pricing with additional interesting topics including inventory constraints, fairness, privacy protection and more. We refer the readers to some recent works that could potentially incorporate our method in future research. In terms of joint pricing and inventory control problem, Chen et al. (2021) develop learning algorithm with censored demand and approximates $f(d_{t},p_{t},y_{t})$ nonparametrically, where $y_{t}$ represents inventory levels. Feng and Zhu (2023) further explore the impact of price protection guarantee, where the customers who purchased a product are endowed the right to receive a refund if the seller lowers the price during the price protection period. Li and Zheng (2023) take the inventory constraints into consideration where the inventory level B is given at the beginning and cannot be replenished. All those studies develop learning algorithms to estimate the underlying unknown demand functions, where our method could be seamlessly integrated. In terms of fairness, Maestre et al. (2018) consider the fairness in price, they introduce Jain’s index to measures the fairness in terms of the average price across groups, while Cohen et al. (2021) estimate unknown demand function under both price fairness and time fairness.
@@
Additional constraints or addition conditions of related interest in dynamic pricing are explored in the following works. Chen et al. (2021) adopt the fundamental framework of differential privacy when implementing demand learning, and develop privacy-preserving dynamic pricing, which tries to maximize the retailer revenue while avoiding leakage of individual customer’s information. Keskin and Zeevi (2017) adopt a linear underlying demand model but assumes the unknown parameters can change over time. van den Boer and Keskin (2022) take reference-price effects into account and learn demand uncertainty by adding the reference-price updating mechanism with unknown parameters. Bastani et al. (2021) leverage a potentially shared structure of the demand function across multiple products and conduct a meta dynamic pricing algorithm to make pricing more efficiently.
@@ 2 Doubly Non-Parametric Learning and Decision
We propose an Algorithm 1 which outputs the empirical optimal policy for contextual dynamic pricing and balances the exploration-exploitation trade-off. The horizon of $T$ periods is divided into episodes with increasing length. Specifically, the $k$th episode has length $n_{k}=2^{k-1}n_{0}$, where $k\geq 1$ and $n_{0}$ is the minimal episode length. For each $k\geq 1$, the $k$th episode is further divided into exploration and commit (or exploitation) phases, which are denoted by ${\cal T}_{k,exp}$ and ${\cal T}_{k,com}$, respectively. In the exploration phase ${\cal T}_{k,exp}$, we collect sample $(x_{t},y_{t})$ through randomly setting price $p_{t}\stackrel{{\scriptstyle d}}{{\sim}}\operatorname{Unif}(0,B)$ and then estimate both $\mathring{\mu}(\cdot)$ or $F(\cdot)$ nonparametrically using the collected sample $\{(x_{t},y_{t})\}_{t\in{\cal T}_{k,exp}}$. In the exploitation phase ${\cal T}_{k,com}$, we proceed to formulate a data-driven optimal policy leveraging the estimated mean utility and noise distribution learned in the exploration phase. The sizes of the exploration phase of uniform random pricing and the exploitation phase of greedy pricing based on estimation can be determined in theoretical analysis to minimize the total regret. The subsequent subsections introduce our proposed approach for the doubly non-parametric estimation of $\mathring{\mu}(\cdot)$ and $F(\cdot)$.
@@
11      Using samples $\{(\boldsymbol{x}_{t},y_{t})\}_{t\in{\cal T}_{k,exp}}$ to fit $\widehat{\mu}_{k}(\boldsymbol{x}_{t})$ for $t\in{\cal T}_{k,exp}$ according to DNN (9) or TDNN (10).
@@
17             Predict $\widehat{\mu}_{k}(\boldsymbol{x}_{t})$ according to DNN (9) or TDNN (10).